OpenStudy (jazilove):

I WILL GIVE MEDAL Given triangle ABC with A(-4, -2), B(4, 4), and C(18, -8), Write the equation of the line containing the altitude that passes through B in standard form.

OpenStudy (jazilove):

please help

OpenStudy (jazilove):

anyone

OpenStudy (anonymous):

i think for this you'd have to use the distance formula

OpenStudy (jazilove):

how would i set that up

OpenStudy (jazilove):

still very confused

OpenStudy (jazilove):

can you explain more?

OpenStudy (anonymous):

So since the altitude is perpendicular to the opposing side, we have have to find the equation of AC.

OpenStudy (jazilove):

how do i do that

OpenStudy (jazilove):

im sorry im very tired

OpenStudy (anonymous):

that's fine same here

OpenStudy (anonymous):

let's see: \[m = \frac{ -8 - (-2) }{ 18 - (-4) } = \frac{ -6 }{ 22 } = \frac{ -3 }{ 11 }\] But since we're looking for the perpendicular slope, we find the inverse reciprocal: \[\frac{ -3 }{ 11 } \rightarrow \frac{ 11 }{ 3 }\] this is our slope for the altitude. \[y = f(4) = 4 = \frac{ 11 }{ 3 } (4) + b\] now, we just have to pass through point B (4,4). So to get it in standard form: \[y-4 = \frac{ 11 }{ 3 }(x-4)\] \[y - 4 = (\frac{ 11 }{ 3 })x - \frac{ 44 }{ 3 }\] \[(\frac{ -11 }{ 3 })x + y = 4 - \frac{ 44 }{ 3 }\] \[\frac{ -11 }{ 3 }x + y = \frac{ -32 }{ 3 }\] and there you have it. But please double check my work just in case.

OpenStudy (jazilove):

thank you so much

OpenStudy (anonymous):

You're welcome